OpenStudy (anonymous):

The radius of a circle is multiplied by 10. Its area is 100 times bigger than the original circle ?

OpenStudy (lgbasallote):

yep :D

OpenStudy (lgbasallote):

\[\large A = \pi (10r)^2 = 100 (\pi r^2)\]

OpenStudy (anonymous):

thank u LG

OpenStudy (lgbasallote):

no probs :D